OpenStudy (anonymous):

Analytic Trigonometry (again!) (Sin2x + Cos2x)^2 = 1

zepdrix (zepdrix):

Expand the binomial :)\[\Large\rm (a+b)^2=a^2+2ab+b^2\]Yes?

OpenStudy (anonymous):

alright that gives \[\sin2x^2 + 2\sin2xcos2x + \cos2x^2 =1\]

zepdrix (zepdrix):

Ok let's clean it up a little bit so it's easier to read, \[\Large\rm \sin^2(2x) + 2\sin(2x)\cos(2x) + \cos^2(2x) =1\]

zepdrix (zepdrix):

If we rearrange our addition:\[\Large\rm \color{orangered}{\sin^2(2x) +\cos^2(2x)} + 2\sin(2x)\cos(2x) =1\]Do you see anything interesting going on?

OpenStudy (anonymous):

is that an identity I see?

zepdrix (zepdrix):

It should be :) hehe

zepdrix (zepdrix):

Remember your most basic Pythagorean Identity involving sine and cosine?

zepdrix (zepdrix):

I think it's the 2x that is confusing you

OpenStudy (anonymous):

Yes, that equals 1 correct?

zepdrix (zepdrix):

\[\Large\rm \color{orangered}{1} + 2\sin(2x)\cos(2x) =1\]Mmmm ok good!

OpenStudy (anonymous):

Okay. Should we subtract 1 to set 2sin(2x)cos(2x) = 0 ?

zepdrix (zepdrix):

Yes. And further, divide both sides by 2, sin(2x) cos(2x) = 0

OpenStudy (anonymous):

Alright

zepdrix (zepdrix):

Then you apply your Zero-Factor Property again, yes? :) OR OR OR, you can apply your Sine Double Angle if you want to simplify things a bit.

OpenStudy (anonymous):

can we do both the sin and cos double angle properties?

zepdrix (zepdrix):

Mmm don't need the Cosine Double Angle.\[\Large\rm \sin(4x)=2\sin(2x)\cos(2x)\]

zepdrix (zepdrix):

Giving us, sin(4x)=0 yah? :o

OpenStudy (anonymous):

I am not aware of that

OpenStudy (anonymous):

how would I go about solving a 4x then? :)

zepdrix (zepdrix):

This is the identity you know, yes? \[\Large\rm \sin(2u)=2\sin(u)\cos(u)\]Replace all of the u's with 2x.

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

So then your sine function is zero at ummmm 0 and pi, yes?\[\Large\rm \sin(4x)=0\qquad\to\qquad 4x=0+k \pi\]Where k is any integer. We solve for x by dividing both sides by 4.

OpenStudy (anonymous):

This is very new to me xD

zepdrix (zepdrix):

Are you not familiar with the k stuff? Are you only looking for solutions between 0 and 2pi maybe?

zepdrix (zepdrix):

Sine and Cosine are periodic, they will repeat their pattern over and over. So where ever we have a solution, we'll have infinitely maybe cause of the repeating nature of sine.

OpenStudy (anonymous):

Oh, adding from the interval [0, 2pi) would have helped, wouldn't it?

zepdrix (zepdrix):

Yah that helps simplify things a bit :)

OpenStudy (anonymous):

I apologize

zepdrix (zepdrix):

Oh oh, ok so things get a little tricky here.

zepdrix (zepdrix):

We have sine giving us zero at 0, pi, 2pi, 3pi, 4pi, 5pi, and so on.

zepdrix (zepdrix):

So those are the angles for our 4x. But we're interested in x. And we want all of the angles that fall within 2pi.

OpenStudy (anonymous):

yee

zepdrix (zepdrix):

So if we divide each of those angles by 4 we get our values for x.\[\Large\rm 0,~\frac{\pi}{4},~\frac{2\pi}{4},~\frac{3\pi}{4},~\frac{4\pi}{4},~\frac{5\pi}{4},~\frac{6\pi}{4},~\frac{7\pi}{4},~\frac{8\pi}{4}\]

zepdrix (zepdrix):

Ooo I'm getting a different answer set when I graph it... hold up lemme make sure I didn't make a boo boo somewhere :(

OpenStudy (anonymous):

haha okay

zepdrix (zepdrix):

no no we're good :)`

zepdrix (zepdrix):

So I'm taking all of these multiples of pi, which are equivalent to our 4x. Dividing them all by 4 gives us the values for x. Should we exclude any of those based on our interval that they gave us? Any of them larger or equal to 2pi?

OpenStudy (anonymous):

nah

zepdrix (zepdrix):

Hmmm 8pi/4. I think that one is a problem for us. That simplifies to 2pi, yah? :o

OpenStudy (anonymous):

yeh

zepdrix (zepdrix):

lol

OpenStudy (anonymous):

o dang

zepdrix (zepdrix):

Hmm this is a tough problem, 8 solutions D: Your brain esplode already?

OpenStudy (anonymous):

like 7 times already

zepdrix (zepdrix):

Mmm yah trig will do that to ya bruh >.<

OpenStudy (anonymous):

bbbbbrrrruuuuuhhhh

OpenStudy (anonymous):

lol

zepdrix (zepdrix):

Do you understand this step?\[\Large\rm \sin(\theta)=0\]Like would you be able to tell me values for \(\Large\rm \theta\) ?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Because our \(\Large\rm \theta\) is our \(\Large\rm 4x\) in this problem.

OpenStudy (anonymous):

okay

zepdrix (zepdrix):

Ok ok chew on that for a while >.< I think your brain needs some rest :3 mine too..

OpenStudy (anonymous):

haha alright, I'll start on the next problem or do some comp sci